OpenStudy (anonymous):

A professor must randomly select 4 students to participate in a mock debate. There are 20 students in his class. In how many different ways can these students be selected, if the order of selection does not matter? A. 3,945 B. 4,845 C. 2,880 D. 116,280

OpenStudy (solomonzelman):

This time it sounds like it is combinations rather than permutations. \(\normalsize\color{blue}{ 20~C~4 }\)

OpenStudy (solomonzelman):

\(\large\color{blue}{ 20~C~4~=~\frac{20!}{4!(20-4)!} }\)

OpenStudy (solomonzelman):

can you finish the prob?

OpenStudy (anonymous):

i forgot what ! means by a number haha

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 5!= 1 \times 2\times 3 \times 4 \times 5 }\) OR \(\normalsize\color{blue}{ n!= 1 \times 2\times 3 \times 4 \times...\times ~(n-2) \times ~(n-1)~ \times n. }\) and/or See ?

OpenStudy (anonymous):

yes i sees

OpenStudy (anonymous):

can i just have the answer cuz i don't wanna do the solving stuff

OpenStudy (anonymous):

400/400*256 @SolomonZelman

OpenStudy (solomonzelman):

you will have to do solving at some point, do it before it's too late.